You found your dream house. It will cost you $300000 and you will put down $60000 as a down payment. For the rest you get a 30-year 6.0% mortgage. What will be your monthly mortgage payment in $ (assume no early repayment)?
Answer #1

Total cost of house = $300000

Down payment = $60000

So remaining amount = $240000

So, monthly payment will be calculated on the remaining amount, i.e. $240000

So, PV = $240000

r = 6%

n= 12 months

t = 30 years

So, monthly annuity A can be calculated using following formula

A = PV*(r/n)/(1-(1+r/n)^(-n*t))

So, A = 240000*(0.06/12)/(1-(1+0.06/12)^(-360)) = $1438.92

monthly mortgage payment is $1438.92